how many gallons of a 50% antifreeze solution must be mixed with 90 gallons of 15% antifreeze to get a mixture that is 40% antifreeze
Answer #1
You have 90 gallons of a 15% antifreeze solution. let x = the number of gallons of 50% antifreeze solution {{{90*0.15 + 0.5x = 0.40(90+x)}}} Solve for x. {{{13.5 + 0.5x = 36+0.4x}}} {{{0.1x = 22.5}}} x = 225 gallons
answered by: daniyal